Good day,

For a different reaction, Kc= 7.45×106, kf=6.05×104s−1 and kr=8.12×10−3 s−1
. Adding a catalyst increases the forward rate constant to 2.93×107 s−1
. What is the new value of the reverse reaction constant, after adding catalyst?

Thanks
1
Expert's answer
2020-08-11T04:37:30-0400

The equilibrium constant of any reversible reaction system is the ratio of forward and reverse rate constants:

Kc = kf/kr

Adding a catalyst speeds up the forward and reverse reactions equally, so it has no effect on the equilibrium constant:

Kc = 7.45×106

kf = 2.93×107 s−1

kr = kf/Kc

kr = 2.93×107 / 7.45×106 = 3.93 s−1

Answer: 3.93 s−1